OpenStudy (anonymous):

-x + 5 = 12 x = __________ -3(-x + 4) = -9 x = __________ 2(4x + 3) - 9 = 45 x = __________

OpenStudy (anonymous):

1. x=-7

OpenStudy (anonymous):

1. -x + 5 = 12 Subtract 5 from each side. - 5 -5 ----------------- x = -7 2. -3(-x + 4) = -9 Distribute the -3 3x + (-12) = -9 Now you add 12 to both sides. +12 +12 ------------------- 3x = 3 Now divide by 3. /3 /3 ---------- x = 1

OpenStudy (anonymous):

1st- add 5 to both sides then divide by -1 2nd- distribute first, add 12 to both side then divide by 3 3rd-distribute first then subtract 6-9, after that add 3 to both side and divide by 8

OpenStudy (anonymous):

2(4x + 3) - 9 = 45 Distribute the 2 8x + 6 - 9 = 45 Then 6-9 = -3 8x - 3 = 45 Now add 3 to both sides. +3 +3 --------------- 8x = 48 dVidie by 8 /8 /8 ----------- x = 6

OpenStudy (anonymous):

Well in 2nd question it won't be essential to distribute.....

OpenStudy (anonymous):

well you could do it another way but distributing is what I was taught and I find it easier. Yea it might be a lot of work but its get the answer right step by step. Short-cuts are great but it might not always give the right solution to the problem. This is just my way of solving problems. I know you have your own too. So I respect your opinion.